Grade 12Algebra

Find
76.5 ÷ 0.15
Find
76.5 ÷ 0.15
Find
76.5 ÷ 0.15
Find
76.5 ÷ 0.15

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (765/10) ÷ (15/100)
= (765/10) × (100/15)
= (765 × 100)/ (10 × 15)
= (51 × 10)/ (1 × 1)
= 510

Thanks